OpenStudy (anonymous):

I need to use four 7's to get numbers 1-20... HELP!

OpenStudy (osanseviero):

1=7/7 +7-7 2=7/7+7/7 3=(7+7+7)/7 4=77/7-7 5=7-(7+7)/7 6=√(7·7)-7/7 7=√(7·7)+7-7 8=√(7·7)+7/7 9=7+(7+7)/7 10=(77-7)/7 11=7/.7+7/7 12=(77+7)/7 13=7+7-7/7 14=(7+7)·(7/7) 15=7+7+7/7 16=7/(.7periodic)+√(7·7) 17=7/.7+raiz(7·7) 18=77/7+7 19=7/(.7periodic)+7/.7 20=7/.7+7/.7

OpenStudy (osanseviero):

http://answers.yahoo.com/question/index?qid=20071207133255AAS64nr (sorry the format here didnt pasted well)